OpenStudy (anonymous):

If a rectangle measures 16 inches by 1 foot, whats its area??

OpenStudy (anonymous):

16inch*ft or 16inch*12inch=(16*12)inch^2 192inch^2

OpenStudy (anonymous):

or 16/12ft*1ft= 16/12ft^2= 4/3ft^2

OpenStudy (anonymous):

sooooo A=192 feet????

OpenStudy (anonymous):

192 inches^2 or 4/3ft^2